OpenStudy (anonymous):

write an equation for the translation of y=6/x that has the asymptotes x=4 and y=5

OpenStudy (anonymous):

and this one: What are the points of discontinuity? y=x-2/x^2+5x-6

zepdrix (zepdrix):

Do you know where the asymptotes are for this function? \[\Large\rm y=\frac{6}{x}\] What? At y=0 and x=0 you say? Ah yes very good! Lemme show you how I like to deal with this type of problem. It might not be how your teacher explained it, but it's worth checking out at least. :) To shift to our asymptotes, we `subtract` the new value from our x or y. So to shift our asymptote 4 to the `right` (moving it to x=4), we subtract 4 from x.\[\Large\rm y=\frac{6}{x-4}\]Then to shift it `up` 5 (moving it to y=5), we subtract 5 from y,\[\Large\rm y-5=\frac{6}{x-4}\]But they will want this in the form y= so we need to add 5 to each side to get it into that form.

OpenStudy (anonymous):

Thankyou so much :)

OpenStudy (anonymous):

Would it be Y= 6/x-4+5?

zepdrix (zepdrix):

Yes, but be careful the way you write that in text! I think what you meant to write was, y=6/(x-4)+5 <- this shows that both the x and 4 are in the denominator.

zepdrix (zepdrix):

`What are the points of discontinuity?`\[\Large\rm y=\frac{x-2}{x^2+5x-6}\]

zepdrix (zepdrix):

To do this one, you need to start by `factoring` the denominator. Understand how to do that?

OpenStudy (anonymous):

No. :(

zepdrix (zepdrix):

\[\Large\rm x^2+5x-6\]So you need two numbers that: `multiply to -6` `add to 5` Let's try ummm,\[\Large\rm -2\cdot3=-6\]\[\Large\rm \color{red}{-2+3\ne5}\]Mmmm -2 and 3 didn't work out :( Can you think of any other factors of -6?

OpenStudy (anonymous):

-3? 1?

zepdrix (zepdrix):

\[\Large\rm \color{red}{-3\cdot1\ne -6}\]Hmm nope those are not factors of -6 :[